# CAGE: Causal Attention Enables Data-Efficient Generalizable Robotic Manipulation

### v1

@@ II-A Generalizable Robotic Manipulation
Current generalist robotic manipulation policies [mt-act, rt1, openvla, oxe, peract, octo, gnfactor, same, rt2] are mainly trained under a multi-task setting using extensive robot demonstration data [mt-act, rt1, oxe, peract], and generalization in these policies is predominantly discussed in the context of unseen tasks. Given the multi-task training setup and large volumes of data, it can be straightforward to find similar task combinations within the training data that may resemble the “unseen” tasks. As a result, as previously discussed in §I, it is challenging to determine whether the policy truly generalizes.
@@ II-B Visual Representation Learning for Manipulation
Inspired by recent advances in parameter-efficient fine-tuning of foundation models, we apply LoRA [lora] for fine-tuning the VFM (DINOv2 [dino]) to enhance observation perception effectively. Perhaps the most similar to our work is SAM-E [same], which also leverages the VFM (SAM [sam]) with LoRA fine-tuning for image perception. Nonetheless, as demonstrated in §IV-C, naively substituting visual encoders with VFMs yields limited generalization improvements, underscoring the effectiveness of our entire policy architecture.
@@ III-B Causal Perceiver for Token Compression
Figure 1: Tasks with Generalization Variations. (a) Transport: the robot needs to localize the object on the left side of the workspace (reach), grasp it (grasp), and drop it to the right side (drop). (b) Dumping: the robot needs to grasp the cup horizontally (grasp), move it over the bowl, tilt it until all balls are out into the bowl (pour), and place the cup in the target area (place). (c) Chopping: the robot needs to grasp the kitchen knife (grasp), use the knife to chop the object [^2] We use real peppers during data collection. However, to avoid unnecessary waste, we use clay to make a pepper for evaluation.on the cutting board into several pieces (chop), and place the knife safely on the foam pad (place). Generalization variations include different backgrounds, objects, and camera views (please refer to §IV-A6 for detailed explanations).
@@ IV Experiments
TABLE II: Results of the $L_{2}$ Evaluation. The $L_{2}$ evaluation is conducted in unseen environments, with different backgrounds, camera views and unseen objects. CAGE largely outperforms all baselines, showcasing a strong generalization ability to out-of-distribution environments.
@@
TABLE III: Results of the $L_{1}$ Evaluation. The $L_{1}$ evaluation is conducted in similar environments, with one of the variations in backgrounds, objects and camera views. The overall success rates are reported for the Transport task here. CAGE shows effective generalization across various similar environments, with acceptable performance degradation.
@@ IV-A1 Platform
We employ a Flexiv Rizon robotic arm [flexiv] with a Dahuan AG-95 gripper [dahuan] as our robot platform. A rectangle area of $45\text{cm}\times 60\text{cm}$ in the center of the tabletop is selected as the robot workspace. We adopt a binocular vision setting with 2 Intel RealSense D435 cameras [realsense]: one global camera in front of the robot facing towards the workspace, and one in-hand camera above the gripper on the end of the robotic arm.
@@ IV-A2 Baselines
We choose the 2D image-based Diffusion Policy (DP) [dp], and the current state-of-the-art 3D point-cloud-based policy RISE [rise] with decent generalization performance in the real world as baselines. For DP, RGB images from both cameras are used, whereas the RGB-D frames from only the global camera are provided to RISE following its original implementation. Both methods are trained from scratch with their recommended hyper-parameter settings. Please refer to Appendix -A for details.
@@ IV-A5 Evaluation Protocols
Evaluations are conducted on a workstation with an Intel Core i9-10900K CPU and an NVIDIA RTX 3090 GPU linked to the robot platform. During the evaluation, we run the inference process and the execution process in parallel for continuous control with a frequency of 10Hz. During the evaluations, we randomly and uniformly generate 15 object positions beforehand for each task and initialize the workspace according to the object positions, ensuring a fair, consistent, and reproducible comparison across different methods. We report stage-wise success rates (the stages are defined in the caption of Fig. 2) to evaluate the performance throughout the task. For the Dumping and the Chopping tasks, we also report the task completion rates, including the average number of balls poured into the bowl (# ball), the average number of poured balls if successfully grasped (# ball (g)), and the average number of chops (# chop) respectively.
@@ IV-A6 Generalization Levels
  • L1 generalization denotes a similar environment with the variation in one of the three aforementioned aspects, denoted as $L_{1,bg}$, $L_{1,obj}$ and $L_{1,cam}$ respectively;
  • L2 generalization allows variations on all three aspects, resulting in an unseen environment for the policy.
@@ IV-A7 Implementations
For each task, CAGE is trained with a batch size of 64 on 4 NVIDIA A100 80GB GPUs for 500 epochs. We set $N_{c}=2$ based on the binocular vision settings in our robot platform. For the vision encoder, we choose DINOv2-large [dino] as the backbone and set the LoRA [lora] rank to 16. We set the observation horizon $T_{o}=4$ and the action horizon $T_{a}=20$. Unlike prior works [dp, umi, act], the data preprocessings are limited to color jitter and center crop to ensure a out-of-distribution evaluation of camera viewpoint generalization. However, we conduct experiments with additional geometric augmentations to show the full potential of CAGE in §IV-C.
Ablation $L_{0}$ $L_{1,bg}$ $L_{1,obj}$ $L_{1,cam}$ Inference Speed
reach grasp drop reach grasp drop reach grasp drop reach grasp drop
CAGE (ours) 1.00 1.00 1.00 0.87 0.87 0.87 0.93 0.80 0.80 1.00 0.80 0.80 280 ms
- Attn-UNet 0.80 0.80 0.80 0.80 0.67 0.67 0.73 0.73 0.73 0.80 0.67 0.67 175 ms
- Causal Obs. Perceiver 0.47 0.40 0.40 0.40 0.13 0.13 0.27 0.00 0.00 0.40 0.20 0.20 160 ms
- DINOv2333After removing DINOv2, the remaining model is still not identical to DP [dp], as we opt for an observation horizon of 4 rather than 2 and use proprioception as the prefix to the noisy actions. Please refer to the appendix for detailed explanations. 0.93 0.73 0.73 0.00 0.00 0.00 0.00 0.00 0.00 0.67 0.53 0.47 100 ms
DP [dp] 0.93 0.73 0.73 0.00 0.00 0.00 0.00 0.00 0.00 0.53 0.27 0.27 100 ms
TABLE IV: Ablation Results in Transport Task. Three key components are gradually removed from CAGE.
@@ IV-B3 Unseen Environment
IV-B4 Out-of-box Evaluation after RH20T Pre-Training
Furthermore, we train CAGE on the same Transport task from RH20T [rh20t] with 164 demonstrations collected by 14 operators under 4 different embodiments. This dataset contains diverse distractors and rich variations in object sizes, backgrounds, and camera views. We design a blue cuboid differing from the training blocks in texture and shape simultaneously as the target for evaluation, since the test object (football) is already used as a distractor in this dataset. As illustrated in Tab. V, we observe that CAGE is capable of operating out-of-box and achieving comparable performance against being trained with a semi-in-domain high-quality dataset (only same in the camera view). It is also worth noting that the workspace definition of RH20T differs from our settings, showcasing a descent degree of workspace generalizability of CAGE beyond that observed in visual aspects.
@@ IV-B4 Out-of-box Evaluation after RH20T Pre-Training
TABLE V: Results of the Out-of-Box Evaluation. The figure on the left illustrates the out-of-box environments for CAGE pre-trained on RH20T [rh20t]. Note that none of the object, background, and camera views appear in the training dataset.
IV-C Ablations
IV-C1 Model Components
As shown in Tab. IV, we observe significant performance gains across the base and $L_{1}$ levels of evaluations with the introduction of Causal Observation Perceiver (2nd row v.s. 3rd row). This validates our statements in §III that Causal Observation Perceiver can effectively retain both spatial and temporal information, facilitating the extraction of features for manipulation.
@@ IV-C1 Model Components
IV-C2 Geometric Augmentation
In our initial setup (§IV-A7), we intentionally exclude random cropping and other augmentation methods, creating a mono-distributed training environment. We now unlock the full potential of CAGE by applying geometric augmentations, including random perspective transform followed by random cropping. The results shown in Tab. VI demonstrate a significant boost in generalization abilities, notably with an 82.5% increase in success rate in the $L_{2}$ evaluation. Therefore, for better performance in the challenging $L_{2}$ evaluations, we include these geometric augmentations for all image-based policies.
@@ IV-C2 Geometric Augmentation
TABLE VI: Ablation Results w/wo Geometric Augmentations.
@@
TABLE VII: Ablation Results of the Observation Horizon.
IV-C3 Observation Horizon
Unlike DP [dp], CAGE can handle long observation horizon from multiple cameras efficiently. We show the performance comparisons of the observation horizon $T_{o}$ in Tab. VII, and find that $T_{o}=4$ performs better than the original DP implementation ($T_{o}=2$), demonstrating the superiority of our approach in effectively extracting and utilizing temporal information from a larger history window.
V Conclusion
Acknowledgement
We would like to thank Chenxi Wang for the insightful discussions, Yiming Wang and Zihao He for their help during the data collection process.
@@ Acknowledgement
We provide detailed information about the architecture of CAGE and our implementations of the selected baselines DP [dp] and RISE [rise]. Main hyper-parameters used for training and evaluation are presented in Tab. VIII for comparison.
@@
Our model uses DINOv2-large [dino] as the vision encoder backbone. In order to preserve its rich semantic knowledge learned from pre-training, the weights of DINOv2 are kept frozen and we employ LoRA [lora] on the projector matrices of the attention modules for parameter-efficient fine-tuning. LoRA rank is set to 16 and a dropout rate of 0.1 is applied to avoid over-fitting. We set the observation horizon $T_{o}=4$ based on the ablation results (Tab. VII in §IV-C). The images are first resized to $256\times 256$ and then cropped to $224\times 224$ as observation inputs. The vision encoder outputs $L=256$ tokens of dimension $D=1024$ for each frame. These tokens are first projected into $D_{\text{token}}=512$. Then, $N_{c}\times T_{o}\times L=2\times 4\times 256=2048$ tokens are compressed into $T_{o}=4$ tokens via our proposed Causal Observation Perceiver. The perceiver is composed of a cross-attention module followed by self-attention modules with causal masks: $T_{o}$ learnable tokens are cross-attended to concatenated observation tokens using a causal mask for each camera; the output tokens are then passed into a sequence of $N=4$ causal self-attention modules for further processing. Dropout layers with $p=0.1$ are also applied on each attention block for a more robust feature extraction. Modified from the pure CNN-based noise prediction network [dp], we append a Cross-Attention Block to each Residual Block for separate conditioning of observation and timestep. We adopt the default setting of a 3-layer UNet with increased channel dimensions 256-512-1024 as suggested in [dp]. Only the color jitter is included as the data augmentation for $L_{0}$ and $L_{1}$ generalization evaluations, and additional geometric augmentation (§IV-C) is applied in $L_{2}$ evaluations for better overall performance.
@@
The CNN-based Diffusion Policy is implemented following the official implementation, except that the vision encoder is upgraded from ResNet-18 to ResNet-50 for a higher capacity in visual expressiveness. Given that DP performs better with a shorter observation horizon, we keep $T_{o}=2$ and $T_{a}=16$. All other hyper-parameters are aligned with those of CAGE to facilitate direct comparisons. The augmentations also follows the CAGE implementations. Moreover, a strictly aligned $T_{o}=4$ and $T_{a}=20$ version of DP is included in Tab. IV (4th row) for the ablation study.
@@
The configuration of RISE remains the same as the official implementation, except that we enable the color jitter augmentation option during training for fair generalization comparisons. Following their original settings, RISE only takes the point cloud from a single global camera with observation horizon $T_{o}=1$ as inputs, and outputs continuous actions for $T_{a}=20$ steps. RISE adopts absolute action representations in the global camera coordinates.
@@
Action Representation relative relative absolute
Rotation Representation 6D [rot6d] 6D [rot6d] 6D [rot6d]
@@
Camera Setting 1 global + 1 in-hand 1 global + 1 in-hand 1 global
@@
TABLE VIII: Detailed Comparison of All Policies.
@@
Figure 2: Observation Images from Different Cameras. The front camera is used as the global camera for training and camera-in-domain experiments ($L_{0}$, $L_{1,bg}$ and $L_{1,obj}$). The upper-left and lower-right cameras are utilized for camera-out-of-domain generalization experiments ($L_{1,cam}$ and $L_{2}$). Both alternative viewpoints introduce a considerable amount of visual distortion.
@@
In order to evaluate the generalization performance of each policy, we devise different extents of distribution shifts on three principal visual factors that are particularly susceptible to variation upon deployment in the real world.
  • Background: The original green background is changed with a white leather-textured one or a tablecloth with light and dark geometric patterns.
  • Object: We modify the target object’s color and geometry to evaluate the object-level generalizability of manipulation policies. For the Dumping and the Chopping tasks, in addition to the change of the target object (bowl in Dumping and pepper in Chopping), we also alter the color of the cardboard (in Dumping) or the foam pad (in Chopping). In the Transport task, we additionally test the policies using a football, which poses a serious challenge due to the large difference in geometry and the need for strong self-correction skills following unsuccessful attempts.
  • Camera View: We introduce significant visual distortions by displacing the camera approximately 16cm away from the original position. Additionally, the original global Intel RealSense D435 camera [realsense] is replaced with an Intel RealSense D415 camera [realsense-415] for a more challenging setup. The images from different cameras are shown in Fig. 2, from which we can observe a significant view discrepancy between training and evaluation cameras.
Thus, we define three levels of generalization based on the aforementioned three aspects.
  • L0 generalization refers to evaluations in the training environments. None of the visual aspects have been changed.
  • L1 generalization denotes a similar environment with the variation in one of the previous three aspects, which we refer to as $L_{1,bg}$, $L_{1,obj}$ and $L_{1,cam}$ respectively;
  • L2 generalization poses variations on all three aspects, resulting in an unseen environment for the policy.
Moreover, we use a script [^4] The script for generating random initial object positions is also provided in our code repository. to generate randomly and uniformly distributed evaluation positions beforehand for each task during evaluations, ensuring a fair, consistent, and reproducible comparison across different models under the same test, as shown in Fig. 3 for the example of the Transport task.
Figure 3: Evaluation Initialization for the Transport Task. The block position is randomly and uniformly initialized before evaluations. These positions remain consistent among evaluations for different policies to ensure comparable and reproducible results.

### v2

@@ II-A Generalizable Robotic Manipulation
Current generalist robotic manipulation policies [mt-act, rt1, oxe, openvla, peract, octo, gnfactor, same, rt2] are mainly trained under a multi-task setting using extensive robot demonstration data [mt-act, rt1, oxe, peract], and generalization in these policies is predominantly discussed in the context of unseen tasks. Given the multi-task training setup and large volumes of data, it can be straightforward to find similar task combinations within the training data that may resemble the “unseen” tasks. As a result, as previously discussed in §I, it is challenging to determine whether the policy truly generalizes.
@@ II-B Visual Representation Learning for Manipulation
Inspired by recent advances in parameter-efficient fine-tuning of foundation models, we apply LoRA [lora] for fine-tuning the VFM (DINOv2 [dino]) to enhance observation perception effectively. Perhaps the most similar to our work is SAM-E [same], which also leverages the VFM (SAM [sam]) with LoRA fine-tuning for image perception. Nonetheless, as demonstrated in §IV-E, naively substituting visual encoders with VFMs yields limited generalization improvements, underscoring the effectiveness of our entire policy architecture.
@@ III-B Causal Perceiver for Token Compression
Figure 1: Tasks with Generalization Variations. (a) Transport: the robot needs to localize the object on the left side of the workspace (reach), grasp it (grasp), and drop it to the right side (drop). (b) Dumping: the robot needs to grasp the cup horizontally (grasp), move it over the bowl, tilt it until all balls are out into the bowl (pour), and place the cup in the target area (place). (c) Chopping: the robot needs to grasp the kitchen knife (grasp), use the knife to chop the object [^2] We use real peppers during data collection. However, to avoid unnecessary waste, we make a pepper out of clay for evaluation.on the cutting board into several pieces (chop), and place the knife safely on the foam pad (place). Generalization variations include different backgrounds, objects, and camera views (please refer to §IV-A6 for detailed explanations).
@@ IV Experiments
TABLE II: Results of the $L_{2}$ Evaluation. The $L_{2}$ evaluation is conducted in unseen environments, with different backgrounds, camera views and unseen objects. CAGE largely outperforms all baselines in all tasks.
@@
TABLE III: Results of the $L_{1}$ Evaluation. The $L_{1}$ evaluation is conducted in similar environments, with one of the variations in backgrounds, objects and camera views. The overall success rates are reported for the Transport task here. CAGE surpasses RISE on success rates and completion rates across all tasks.
@@ IV-A1 Platform
We employ a Flexiv Rizon robotic arm [flexiv] with a Dahuan AG-95 gripper [dahuan] as our robot platform. A rectangle area of $45\text{cm}\times 60\text{cm}$ in the center of the tabletop is selected as the robot workspace. We adopt a binocular vision setting with 2 Intel RealSense D435 cameras [realsense]: one fixed camera in front of the robot facing towards the workspace, and one in-hand camera above the gripper on the end of the robotic arm.
@@ IV-A2 Baselines
We choose the 2D image-based Diffusion Policy (DP) [dp], and the current state-of-the-art 3D point-cloud-based policy RISE [rise] with decent generalization performance in the real world as baselines. For DP, RGB images from both cameras are used, whereas the RGB-D frames from only the fixed camera are provided to RISE following its original implementation. Both methods are trained from scratch with their recommended hyper-parameter settings. Please refer to Appendix -A for details.
@@ IV-A5 Evaluation Protocols
Evaluations are conducted on a workstation with an Intel Core i9-10900K CPU and an NVIDIA RTX 3090 GPU linked to the robot platform. During the evaluation, we run the inference process and the execution process in parallel for continuous control with a frequency of 10Hz. During the evaluations, we randomly and uniformly generate test positions beforehand for each task and initialize the workspace accordingly, ensuring a fair, consistent, and reproducible comparison across different methods. We report stage-wise success rates (the stages are defined in the caption of Fig. 2) to evaluate the performance throughout the task. For the Dumping and the Chopping tasks, we also report the task completion rates, including the average number of balls poured into the bowl (# ball), the average number of poured balls if successfully grasped (# ball (g)), and the average number of chops (# chop) respectively.
@@ IV-A6 Generalization Levels
  • L1 generalization refers to evaluations in a similar environment with the variation in one of the three aforementioned aspects, denoted as $L_{1,bg}$, $L_{1,obj}$ and $L_{1,cam}$ respectively;
  • L2 generalization refers to evaluations in an unseen environment with variations present across all three aspects, thereby emulating the scenario in real-world deployment.
@@ IV-A7 Implementations
For each task, CAGE is trained with a batch size of 64 on 4 NVIDIA A100 80GB GPUs for 500 epochs. We set $N_{c}=2$ based on the binocular vision settings in our robot platform. For the vision encoder, we choose DINOv2-large [dino] as the backbone and set the LoRA [lora] rank to 16. We set the observation horizon $T_{o}=4$ and the action horizon $T_{a}=20$ with a relative action representation [umi]. Unlike prior works [dp, umi, act], the data preprocessings are limited to color jitter and center crop to ensure a out-of-distribution evaluation of camera viewpoint generalization. However, we conduct experiments with additional geometric augmentations to show the full potential of CAGE in §IV-E.
@@ IV-B Results
Platform Transport, similar Dumping, similar
Robot Gripper reach grasp drop grasp pour place # ball # ball (g)
Flexiv Dahuan 1.00 1.00 1.00 0.80 0.80 0.50 7.5 / 10 9.4 / 10
Flexiv Robotiq 0.73 0.60 0.60 0.40 0.40 0.40 3.2 / 10 8.0 / 10
RealMan Dahuan 0.53 0.47 0.47 0.20 0.20 0.10 1.2 / 10 6.0 / 10
TABLE IV: Cross-Embodiment Evaluation Results under Similar Environment. Similar denotes a similar environment setup with a green background, the same objects, and a close camera view. The background and the camera are not identical to the training environment as their platforms are different. New hardware is highlighted in yellow.
Platform Transport, novel Dumping, novel
Robot Gripper reach grasp drop grasp pour place # ball # ball (g)
Flexiv Dahuan 0.93 0.73 0.73 0.80 0.60 0.20 3.4 / 10 4.3 / 10
Flexiv Robotiq 0.60 0.53 0.53 0.40 0.40 0.40 3.2 / 10 8.0 / 10
RealMan Dahuan 0.33 0.13 0.13 0.10 0.00 0.00 0.0 / 10 0.0 / 10
TABLE V: Cross-Embodiment Evaluation Results under Novel Environment. We further replace the background and the target objects following the same protocol as described in §IV-A6. New hardware is highlighted in yellow.
@@ IV-B3 Unseen Environment
IV-C Cross-Embodiment Evaluations
Apart from visual generalization capabilities, it is essential for a generalizable policy to effectively perform the learned tasks on new embodiments. Despite the inherent challenges, we report the task success rates for the Transport and Dumping tasks, as well as the completion rates for the latter, on new platforms with different hardware configurations.
The evaluation results on two similar platforms featuring either a new gripper or a different type of robotic arm are presented in Tab. IV and Tab. V. A significant performance drop is expected given the entirely new workspaces and manipulation interfaces. Nonetheless, for straightforward pick-and-place tasks like Transport, CAGE achieves a near 50% success rate in most cases, except for configurations involving the RealMan arm in a novel environment.
Additionally, we observe that for more delicate tasks, our model demonstrates greater robustness to changes in the gripper compared to variations in the robotic arm. While these results do not yet meet the standard of excellence, we hope they can serve as a valuable baseline for future efforts aimed at enhancing cross-embodiment generalization performance.
IV-D RH20T Pre-training
@@
TABLE VI: Results of the Out-of-Box Evaluation. The figure on the left illustrates the out-of-box environments for CAGE pre-trained on RH20T [rh20t]. Note that none of the object, background, and camera views appear in the training dataset.
In addition to the self-collected, high-quality dataset, we train CAGE on the same Transport task from RH20T [rh20t] to evaluate the out-of-box generalization ability after pre-training. This dataset contains 164 demonstrations, collected by 14 operators under 4 different embodiments, and incorporates diverse distractors and rich variations in object sizes, backgrounds, and camera views. For evaluation, we design a blue cuboid [^3] The originally purposed test object (football) is already used as a distractor in this dataset. that differs from the training blocks both in texture and shape as the target. As illustrated in Tab. VI, we observe that while baselines fail to perform the task, CAGE is capable of operating out-of-box and achieving comparable performance against being trained with a semi-in-domain high-quality dataset (only same in the camera view). This proves that our approach can be scaled up to large and diverse datasets and yield a corresponding performance improvement.
IV-E Ablations
Ablation $L_{0}$ $L_{1,bg}$ $L_{1,obj}$ $L_{1,cam}$
reach grasp drop reach grasp drop reach grasp drop reach grasp drop
CAGE (ours) 1.00 1.00 1.00 0.87 0.87 0.87 0.93 0.80 0.80 1.00 0.80 0.80
- Attn-UNet 0.80 0.80 0.80 0.80 0.67 0.67 0.73 0.73 0.73 0.80 0.67 0.67
- Causal Obs. Perceiver 0.47 0.40 0.40 0.40 0.13 0.13 0.27 0.00 0.00 0.40 0.20 0.20
- DINOv2444After removing DINOv2, the remaining model is still not identical to DP [dp], as we opt for an observation horizon of 4 rather than 2 and use proprioception as the prefix to the noisy actions. Please refer to Appendix -A for detailed explanations. 0.93 0.73 0.73 0.00 0.00 0.00 0.00 0.00 0.00 0.67 0.53 0.47
DP [dp] 0.93 0.73 0.73 0.00 0.00 0.00 0.00 0.00 0.00 0.53 0.27 0.27
TABLE VII: Ablation Results in Transport Task. Three key components are sequentially removed from CAGE.
IV-E1 Model Components
As shown in Tab. VII, we observe significant performance gains across the base and $L_{1}$ levels of evaluations with the introduction of Causal Observation Perceiver (2nd row v.s. 3rd row). This validates our statements in §III that Causal Observation Perceiver can effectively retain both spatial and temporal information, facilitating the extraction of features for manipulation.
@@ IV-E1 Model Components
IV-E2 Geometric Augmentation
@@
TABLE VIII: Ablation Results w/wo Geometric Augmentations.
In our initial setup (§IV-A7), we intentionally exclude random cropping and other augmentation methods, creating a mono-distributed training environment. We now unlock the full potential of CAGE by applying geometric augmentations, including random perspective transform followed by random cropping. The results shown in Tab. VIII demonstrate a significant boost in generalization abilities, notably with an 82.5% increase in success rate in the $L_{2}$ evaluation. Therefore, for better performance in the challenging $L_{2}$ evaluations, we include these geometric augmentations for all image-based policies.
IV-E3 Observation Horizon
Unlike DP [dp], CAGE can handle long observation horizon from multiple cameras efficiently. We show the performance comparisons of the observation horizon $T_{o}$ in Tab. IX, and find that $T_{o}=4$ performs better than the original DP implementation ($T_{o}=2$), demonstrating the superiority of our approach in effectively extracting and utilizing temporal information from a larger history window.
V Conclusion
Acknowledgement
@@
TABLE IX: Ablation Results of the Observation Horizon.
We would like to thank Chenxi Wang for insightful discussions and for setting up the RealMan platform in cross-embodiment evaluations, Yiming Wang and Zihao He for their help during the data collection process.
@@
We provide detailed information about the architecture of CAGE and our implementations of the selected baselines DP [dp] and RISE [rise]. Main hyper-parameters used for training and evaluation are presented in Tab. X for comparison.
@@
Our model uses DINOv2-large [dino] as the vision encoder backbone. In order to preserve its rich semantic knowledge learned from pre-training, the weights of DINOv2 are kept frozen and we employ LoRA [lora] on the projector matrices of the attention modules for parameter-efficient fine-tuning. LoRA rank is set to 16 and a dropout rate of 0.1 is applied to avoid over-fitting. We set the observation horizon $T_{o}=4$ based on the ablation results (Tab. IX in §IV-E). The images are first resized to $256\times 256$ and then cropped to $224\times 224$ as observation inputs. The vision encoder outputs $L=256$ tokens of dimension $D=1024$ for each frame. These tokens are first projected into $D_{\text{token}}=512$. Then, $N_{c}\times T_{o}\times L=2\times 4\times 256=2048$ tokens are compressed into $T_{o}=4$ tokens via our proposed Causal Observation Perceiver. The perceiver is composed of a cross-attention module followed by self-attention modules with causal masks: $T_{o}$ learnable tokens are cross-attended to concatenated observation tokens using a causal mask for each camera; the output tokens are then passed into a sequence of $N=4$ causal self-attention modules for further processing. Dropout layers with $p=0.1$ are also applied on each attention block for a more robust feature extraction. Modified from the pure CNN-based noise prediction network [dp], we append a Cross-Attention Block to each Residual Block for separate conditioning of observation and timestep. We adopt the default setting of a 3-layer UNet with increased channel dimensions 256-512-1024 as suggested in [dp]. For the data augmentation, only the color jitter is included in $L_{0}$ and $L_{1}$ generalization evaluations, and additional geometric augmentation (§IV-E) is applied in $L_{2}$ evaluations for better overall performance.
@@
The CNN-based Diffusion Policy is implemented following the official implementation, except that the vision encoder is upgraded from ResNet-18 to ResNet-50 for a higher capacity in visual expressiveness. Given that DP performs better with a shorter observation horizon, we keep $T_{o}=2$ and $T_{a}=16$. All other hyper-parameters are aligned with those of CAGE to facilitate direct comparisons. The augmentations also follows the CAGE implementations. Moreover, a strictly aligned $T_{o}=4$ and $T_{a}=20$ version of DP is included in Tab. VII (4th row) for the ablation study.
@@
The configuration of RISE remains the same as the official implementation, except that we enable the color jitter augmentation option during training for fair generalization comparisons. Following their original settings, RISE only takes the point cloud from a single fixed camera with observation horizon $T_{o}=1$ as inputs, and outputs continuous actions for $T_{a}=20$ steps. RISE adopts absolute action representations in the camera coordinate.
@@
Camera Setting 1 fixed + 1 in-hand 1 fixed + 1 in-hand 1 fixed
@@
TABLE X: Detailed Comparison of All Policies.
@@
Figure 2: Observation Images from Different Cameras. The front camera is used as the fixed camera for training and camera-in-domain experiments ($L_{0}$, $L_{1,bg}$ and $L_{1,obj}$). The upper-left and lower-right cameras are utilized for camera-out-of-domain generalization experiments ($L_{1,cam}$ and $L_{2}$). Both alternative viewpoints introduce a considerable amount of visual distortion.
@@
In order to evaluate the generalization performance of each policy, we devise distribution shifts on three principal visual factors that are particularly susceptible to variation upon deployment in the real world:
  • Background: The original green background is changed to a white leather-textured one or a tablecloth with light and dark geometric patterns depending on the task.
  • Object: We modify the target object’s color and geometry to evaluate the object-level generalizability of manipulation policies. For the Dumping and the Chopping tasks, in addition to the change of the target object (bowl in Dumping and pepper in Chopping), we also alter the color of the cardboard (in Dumping) or the foam pad (in Chopping). In the Transport task, we additionally test the policies using a football, which poses a serious challenge due to the huge difference in geometry and the need for strong self-correction skills following unsuccessful attempts.
  • Camera View: We introduce a large visual distortions by displacing the camera approximately 16cm away from the original position. Additionally, the original fixed Intel RealSense D435 camera [realsense] is replaced with an Intel RealSense D415 camera [realsense-415] for a more challenging setup. The images from different cameras are shown in Fig. 2, from which we can observe a significant view discrepancy between training and evaluation cameras.
Therefore, we define three levels of generalization tests:
  • L0 generalization is conducted in the same environment seen during training. We evaluate the most basic ability of the policy to learn the task from demonstrations.
  • L1 generalization is conducted in a similar environment with the variation in one of the three visual aspects, which we refer to as $L_{1,bg}$, $L_{1,obj}$ and $L_{1,cam}$ respectively. We evaluate the robustness of the policy under moderate distribution shifts.
  • L2 generalization is conducted in a totally unseen environment with variations on all three aspects. We evaluate the capability of the policy to complete the learned task in out-of-domain scenes.
Moreover, we use a script [^5] The script for generating random initial object positions is also provided in our code repository. to generate randomly and uniformly distributed evaluation positions beforehand for each task, ensuring a fair, consistent, and reproducible comparison across different models under the same test, as shown in Fig. 3 for the example of the Transport task.
Figure 3: Evaluation Initialization for the Transport Task. The target block is placed at each designated location for each evaluation round.